(A) Mrs. Smith took a random sample of students after winter break and asked if they enjoyed their break. She found that 70% of the 40 students she asked said yes. Estimate and interpret a 90% confidence interval for the proportion of all students who enjoyed their break.

(B) How large of a sample will allow Mrs. Smith to get a margin of error of 2.5% at a 95% confidence level? Remember to SHOW your work with the numbers plugged into the formula and answer in context.
Answer #1

Solution :

A)

Given that,

n = 40

Point estimate = sample proportion = \hat p = 70% = 0.70

1 - \hat p = 1 - 0.70 = 0.30

At 90% confidence level the z is ,

\alpha = 1 - 90% = 1 - 0.90 = 0.10

\alpha / 2 = 0.10 / 2 = 0.05

Z\alpha/2 = Z0.05 = 1.645

Margin of error = E = Z\alpha / 2 * \sqrt ((\hat p * (1 - \hat p )) / n)

= 1.645 * (\sqrt((0.70 * 0.3.0) / 40)

= 0.119

A 90% confidence interval for population proportion p is ,

\hat p - E < p < \hat p + E

0.70 - 0.119 < p < 0.70 + 0.119

0.581 < p < 0.819

(0.581 , 0.819)

(B)

Given that,

\hat p = 0.70

1 - \hat p = 1 - 0.70 = 0.30

margin of error = E = 2.5% = 0.025

At 95% confidence level the z is ,

\alpha = 1 - 95% = 1 - 0.95 = 0.05

\alpha / 2 = 0.05 / 2 = 0.025

Z\alpha/2 = Z0.025 = 1.96

sample size = n = (Z\alpha / 2 / E )2 * \hat p * (1 - \hat p )

= (1.96 / 0.025)2 * 0.70 * 0.30

= 1291

sample size = 1291

The 95% confidence interval for the population proportion p is :